# Bayesian Experimental Design for Model Discrepancy Calibration: An Auto-Differentiable Ensemble Kalman Inversion Approach

### v1

@@ 2.3 AD-EKI for A Differentiable Estimation of KL Divergence
In general, the ensemble-based KL divergence explicitly depends on the distribution of the ensemble at each EKI iteration, i.e., $\bar{\bm{\theta}}_{n}$ and $\Sigma_{n}^{\bm{\theta}\bm{\theta}}$, which further depends on the input data $\mathbf{y}(\mathbf{d})$ and the predictions $\mathbf{g}^{(j)}_{n-1}(\bm{\theta}^{(j)}_{n-1},\mathbf{d})$. Here we present a brief derivation for the detailed form of $\mathrm{d}D_{\text{KL}}^{e}/\mathrm{d}\mathbf{d}$ to facilitate the discussion of computational complexity. More specifically, considering the $K$-th iteration of EKI, the terms in Eq. (7) that are dependent on $\mathbf{d}$ include the ensemble mean $\bar{\bm{\theta}}_{K}$ and the covariance matrix $\Sigma_{K}^{\bm{\theta}\bm{\theta}}$. The initial ensemble parameters, $\bar{\bm{\theta}}_{0}$ and $\Sigma_{0}^{\bm{\theta}\bm{\theta}}$, are pre-specified and independent of the design $\mathbf{d}$. The gradient $\nabla_{\mathbf{d}}D_{\text{KL}}(\mathbf{d})$ of the KL divergence with respect to the design is explicitly given by
$\frac{\mathrm{d}D_{\text{KL}}^{e}}{\mathrm{d}\mathbf{d}}=$ $\frac{\partial D_{\text{KL}}^{e}}{\partial\bar{\bm{\theta}}_{K}}\frac{\mathrm{d}\bar{\bm{\theta}}_{K}}{\mathrm{d}\mathbf{d}}+\frac{\partial D_{\text{KL}}^{e}}{\partial\Sigma^{\bm{\theta}\bm{\theta}}_{K}}:\frac{\mathrm{d}\Sigma^{\bm{\theta}\bm{\theta}}_{K}}{\mathrm{d}\mathbf{d}}$ (9)
@@
where :::: represents the Frobenius product. Detailed derivation of Eq. (9) can be found in A. The gradient terms at the right-hand side of Eq. (9) can be further expanded as
@@
The mapping $\left[\mathbf{y},\mathbf{g}^{(j)}_{n-1}\right]\mapsto D_{\text{KL}}^{e}$ can be effectively handled by auto-differentiation tools such as PyTorch or JAX. The mapping $\mathbf{d}\mapsto\left[\mathbf{y},\mathbf{g}^{(j)}_{n-1}\right]$ involves the data measurement process and the forward model. In this work, we focus on the data measurement locations as the designs $\mathbf{d}$. If a spatially discretized solution of the full field is available, e.g., from the forward model output, interpolation can be performed to obtain the gradient information with respect to $\mathbf{d}$. In contrast, if the data is directly measured from the true field, infinitesimal perturbation analysis (IPA, [90, 54]) methods provide a feasible approximation. In practice, both the interpolation for the forward model output field and the perturbation-based sensitivity estimation in IPA can be directly approximated using the JAX built-in function jax.scipy.ndimage.map_coordinates.
@@
The computational time complexity of auto-differentiating to the map $\mathbf{d}\mapsto D_{\text{KL}}^{e}$ remains of the same order as that of the forward mapping [69] and scales linearly with the total number of EKI iterations $K$ at $\mathcal{O}(K)$, provided that the forward model of different ensemble members in EKI is simulated in parallel. Implementations of JAX loop functions like jax.scan and jax.fori_loop could help minimize the scaling factor as they directly compile an entire loop into a single XLA kernel without repeating Python interpretation. The computational time used for transferring data and commands between Python to XLA can be minimized from $2K$ times to twice. An increase in time complexity is expected if the forward model of different ensemble members is executed sequentially or partially sequentially due to the memory limit.
@@ 2.4 AD-EKI for A Differentiable Estimation of Expected Information Gain
$D_{\text{KL}}^{e}$ represents the actual information gain provided by some measurement data from the true system. In the context of experimental design, those measurement data are often unavailable a priori before an experiment is designed. Therefore, the expected information gain (EIG) is employed to design an experiment based on the modeled system and prior knowledge without requiring the actual data from the experiments. As expressed in Eq. 3, EIG integrates out the dependency on data $\mathbf{y}$ in the KL divergence $D_{\text{KL}}$ by the prior predictive distribution $p(\mathbf{y}|\mathbf{d})$. The differentiation formulae of $D_{\text{KL}}^{e}$ presented in Section 2.3 remains valid, except for that data $\mathbf{y}$ is predicted by the model when EIG is employed as the utility function of BED.
Specifically, the distribution $p(\mathbf{y}|\mathbf{d})$ of the predicted data $\mathbf{y}$ is induced by the prior distribution $p(\bm{\theta})$ of the parameter $\bm{\theta}$, which is propagated through the PDE solution process under a given design $\mathbf{d}$ and subsequently transformed by the observation model with observing noise $\bm{\eta}\sim\mathcal{N}(\mathbf{0},\Gamma)$: $\{p(\bm{\theta}),\mathcal{N}(\mathbf{0},\Gamma)\}\mapsto p(\mathbf{y}|\mathbf{d})$. While directly evaluating this predictive distribution $p(\mathbf{y}|\mathbf{d})$ and calculating $D_{\text{KL}}^{e}(\mathbf{d,\mathbf{y}^{(m)}})$ for each sampled $\mathbf{y}^{(m)}\sim p(\mathbf{y}|\mathbf{d})$ is possible, this sampling process is not differentiable and reparameterization trick [91] is required.
@@
$\mathbb{E}_{\mathbf{y(\mathbf{d})}\sim p(\mathbf{y}|\mathbf{d})}[D_{\text{KL}}^{e}(\mathbf{d},\mathbf{y})]=\mathbb{E}_{\bm{\theta}\sim p(\bm{\theta}),\mathbf{\mathbf{\eta}}\sim p(\eta)}\{D_{\text{KL}}^{e}[\mathbf{d},\mathbf{y}(\bm{\theta},\mathbf{\eta})]\}\approx\frac{1}{M}\sum_{m=1}^{M}D_{\text{KL}}^{e}[\mathbf{d},\mathbf{y}(\bm{\theta}^{m},\mathbf{\eta}^{m})],$ (12)
where each $D_{\text{KL}}^{e}[\mathbf{d},\mathbf{y}(\bm{\theta}^{m},\mathbf{\eta}^{m})]$ is evaluated by Eq. (7) in this work. Consequently, differentiating EIG with respect to design $\mathbf{d}$ can be estimated as:
$\frac{\mathrm{d}}{\mathrm{d}\mathbf{d}}\mathbb{E}_{\mathbf{y(\mathbf{d})}}[D_{\text{KL}}^{e}(\mathbf{d},\mathbf{y})]$ $\approx\frac{\mathrm{d}}{\mathrm{d}\mathbf{d}}\left[\frac{1}{M}\sum_{m=1}^{M}D_{\text{KL}}^{e}[\mathbf{d},\mathbf{y}(\bm{\theta}^{m},\mathbf{\eta}^{m})]\right]$ (13)
$=\frac{1}{M}\sum_{m=1}^{M}\frac{\mathrm{d}}{\mathrm{d}\mathbf{d}}D_{\text{KL}}^{e}[\mathbf{d},\mathbf{y}(\bm{\theta}^{m},\mathbf{\eta}^{m})],$
where the interchange of the order of differentiation and integration in the second equal is valid with mild assumptions that $D_{\text{KL}}^{e}$ and $\partial D_{\text{KL}}^{e}/\partial\mathbf{d}$ are continuous and an integrable function bounds the partial derivative term.
For a single sample of $\mathbf{y}^{m}$, the computational cost of AD-EKI remains nearly unchanged. As the calculation of $D_{\text{KL}}^{e}[\mathbf{d},\mathbf{y}(\bm{\theta}^{m},\mathbf{\eta}^{m})]$ for each sample is independent, the total computational complexity for $M$ samples scales linearly as $\mathcal{O}(MJ)$. It is also worth noting that sequentially performing AD-EKI for each sample does not change the peak memory cost of $\mathcal{O}(J)$. The estimation of EIG in Eq. (12) differs from a nested Monte Carlo method, which is commonly used in BED to approximate the double integral in EIG. The primary distinction between this formulation of Eq. (12) and the nested Monte Carlo method lies in the estimation of $D_{\text{KL}}^{e}[\mathbf{d},\mathbf{y}(\bm{\theta}^{m},\mathbf{\eta}^{m})]$, which is approximated by the ensemble Kalman inversion in this work. A key advantage is that a small ensemble size is often sufficient for ensemble Kalman methods, even if dealing with a high-dimensional problem, e.g., only an order of $10^{2}$ samples are typically required to handle parameter spaces with dimensions as high as $10^{8}$ in numerical weather prediction, with various techniques developed for a robust covariance estimation [92, 93].
@@ 2.5 A Hybrid Framework for BED with Model Discrepancy
The difference between the true system and a physics-based model, known as model discrepancy, is ubiquitous in the modeling and simulation of complex dynamical systems. Existing data-driven methods on the calibration of model discrepancy often rely on empirical choices of the training data. Unlike the unknown parameters of a physics-based model that are often in a low-dimensional vector space, the model discrepancy is in an infinite-dimensional function space [82], for which the approximation approaches (e.g., based on neural networks) can involve unknown coefficients in a high-dimensional vector space, which is challenging for employing standard BED methods to systematically identify optimal training data. In this work, we aim to build on the AD-EKI introduced in Section 2.4 and develop an efficient framework to identify optimal designs that guide active data acquisition for the calibration of the model discrepancy.
The true system in a general form can be written as:
In this work, the design $\mathbf{d}$ corresponds to the spatiotemporal coordinate $(\mathbf{z},t)$. The measurement $\mathbf{y}$ is the observation of variable $\mathbf{u}$ that represents the state of the physical system at the design $\mathbf{d}$, i.e., $\mathbf{y}=\mathbf{u}(\mathbf{d})+\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{\Sigma}_{\bm{\epsilon}})$ represents the measurement noise which is assumed to be Gaussian in this work when observing the state from the true system.
Although performing BED on the joint parameter space $\{\bm{\theta}_{\mathcal{G}},\bm{\theta}_{\textrm{NN}}\}$ is a natural first thought, incorporating network parameters significantly increases the dimensionality of the parameter space, posing a major challenge for full Bayesian methods due to computational intractability. It is worth noting that some recent works [94, 20, 17, 95] have been developed for Bayesian experimental design with high-dimensional unknown parameters, which rely on building an efficient and generalizable parameter-to-observable map and thus are not directly applicable for the modeled system in Eq. (15). As an efficient alternative to full Bayesian methods, ensemble Kalman methods introduce Gaussian assumptions while demonstrating robust performance on nonlinear problems. Therefore, we propose a hybrid strategy that treats physical parameters with a full Bayesian approach, while dealing with neural network parameters via AD-EKI and gradient-based optimization. More specifically, we decompose the BED problem for $\{\bm{\theta}_{\mathcal{G}},\bm{\theta}_{\textrm{NN}}\}$ into two sub-problems. First, we determine the optimal design for the physical parameters $\bm{\theta}_{\mathcal{G}}$ using standard BED techniques and update the belief distribution through Bayesian inference. Next, we determine the optimal experiment design for the calibration of neural network parameters $\bm{\theta}_{\textrm{NN}}$ by maximizing the ensemble-based utility function enabled by AD-EKI in Section 2.4. With the data from the optimal design, we then update $\bm{\theta}_{\textrm{NN}}$ via a standard gradient-based optimization.
@@
$\mathbf{d}^{*}_{\text{NN}}=\operatorname*{arg\,max}_{\mathbf{d}\in\mathcal{D}}\mathbb{E}[D_{\textrm{KL}}^{e}(p(\bm{\theta}_{\text{NN}}|\mathbf{y};\bm{\theta}_{\mathcal{G}}^{*})||p(\bm{\theta}_{\text{NN}}))],$ (19)
where $D_{\textrm{KL}}^{e}$ denotes the KL-divergence estimated by the AD-EKI introduced in Section 2.3. For the first stage of sequential BED, we assume a Gaussian prior distribution of the neural network coefficients. For each subsequent stage, the posterior distribution from the previous stage serves as the prior distribution $p(\bm{\theta}_{\text{NN}})$. With the optimal design $\mathbf{d}^{*}_{\text{NN}}$ being identified for the neural-network-based model discrepancy, the corresponding data $\mathbf{y}_{\text{NN}}$ is then used to update the network coefficients by a standard gradient-based optimization method. In practice, we leverage the entire accumulated dataset $\mathrm{Y}_{\text{NN}}^{(i-1)}=\{\mathbf{y}_{\text{NN}}^{1},\mathbf{y}_{\text{NN}}^{2},\cdots,\mathbf{y}_{\text{NN}}^{(i-1)}\}$ from previous stages in sequential BED, together with the data $\mathbf{y}_{\text{NN}}^{(i)}$ at the stage $i$ for the training of the neural-network-based model discrepancy.

### v2

@@ 2.3 AD-EKI for A Differentiable Estimation of KL Divergence
In general, the ensemble-based KL divergence explicitly depends on the distribution of the ensemble at each EKI iteration, i.e., $\bar{\bm{\theta}}_{n}$ and $\Sigma_{n}^{\bm{\theta}\bm{\theta}}$, which further depends on the input data $\mathbf{y}(\mathbf{d})$ and the predictions $\mathbf{g}^{(j)}_{n-1}(\bm{\theta}^{(j)}_{n-1},\mathbf{d})$. Here we present a brief derivation for the detailed form of $\mathrm{d}\tilde{D}_{\text{KL}}/\mathrm{d}\mathbf{d}$ to facilitate the discussion of computational complexity. Note that all gradient calculations with respect to the design $\mathbf{d}$ are taken on the ensemble-based utility $\tilde{D}_{\mathrm{KL}}(\mathbf{d})$. More specifically, considering the $K$-th iteration of EKI, the terms in Eq. (9) that are dependent on $\mathbf{d}$ include the ensemble mean $\bar{\bm{\theta}}_{K}$ and the covariance matrix $\Sigma_{K}^{\bm{\theta}\bm{\theta}}$. The initial ensemble parameters, $\bar{\bm{\theta}}_{0}$ and $\Sigma_{0}^{\bm{\theta}\bm{\theta}}$, are pre-specified and independent of the design $\mathbf{d}$. The gradient $\nabla_{\mathbf{d}}\tilde{D}_{\text{KL}}(\mathbf{d})$ of the ensemble-based KL divergence with respect to the design is explicitly given by
$\frac{\mathrm{d}\tilde{D}_{\text{KL}}}{\mathrm{d}\mathbf{d}}=$ $\frac{\partial\tilde{D}_{\text{KL}}}{\partial\bar{\bm{\theta}}_{K}}\frac{\mathrm{d}\bar{\bm{\theta}}_{K}}{\mathrm{d}\mathbf{d}}+\frac{\partial\tilde{D}_{\text{KL}}}{\partial\Sigma^{\bm{\theta}\bm{\theta}}_{K}}:\frac{\mathrm{d}\Sigma^{\bm{\theta}\bm{\theta}}_{K}}{\mathrm{d}\mathbf{d}}$ (11)
@@
where :: represents the Frobenius product. Detailed derivation of Eq. (11) can be found in A. The gradient terms at the right-hand side of Eq. (11) can be further expanded as
@@
The mapping $\left[\mathbf{y},\mathbf{g}^{(j)}_{n-1}\right]\mapsto\tilde{D}_{\text{KL}}$ can be effectively handled by auto-differentiation tools such as PyTorch or JAX. The mapping $\mathbf{d}\mapsto\left[\mathbf{y},\mathbf{g}^{(j)}_{n-1}\right$] involves the data measurement process and the forward model. In this work, we focus on the data measurement locations as the designs $\mathbf{d}$. If a spatially discretized solution of the full field is available, e.g., from the forward model output, interpolation can be performed to obtain the gradient information with respect to $\mathbf{d}$. In contrast, if the data is directly measured from the true field, infinitesimal perturbation analysis (IPA, [ho_perturbation_1983, huan_gradient-based_2014]) methods provide a feasible approximation. In practice, both the interpolation for the forward model output field and the perturbation-based sensitivity estimation in IPA can be directly approximated using the JAX built-in function jax.scipy.ndimage.map_coordinates.
@@
The computational time complexity of auto-differentiating to the map $\mathbf{d}\mapsto\tilde{D}_{\text{KL}}$ remains of the same order as that of the forward mapping [chen_autodifferentiable_2022] and scales linearly with the total number of EKI iterations $K$ at $\mathcal{O}(K)$, provided that the forward model of different ensemble members in EKI is simulated in parallel. Implementations of JAX loop functions like jax.scan and jax.fori_loop could help minimize the scaling factor as they directly compile an entire loop into a single XLA kernel without repeating Python interpretation. The computational time used for transferring data and commands between Python to XLA can be minimized from $2K$ times to twice. An increase in time complexity is expected if the forward model of different ensemble members is executed sequentially or partially sequentially due to the memory limit.
@@
While both AD-EnKF and our AD-EKI differentiate through Kalman updates, our work extends the auto-differentiable filtering paradigm from dynamic state estimation to static inverse problems and Bayesian experimental design. AD-EnKF is tailored to sequential data assimilation: it alternates forecast and analysis steps along a time series and differentiates through each assimilation cycle to improve real-time state. In contrast, our AD-EKI addresses static inverse problems and BED by repeatedly applying the EnKF analysis update conditioned on the same observations. For nonlinear forward models, a one-off Gaussian update yields only a local approximation of the parameter posterior [iglesias_ensemble_2013, schillings2017analysis], which is inadequate for global design objectives. EKI overcomes this by iteratively refining the ensemble through multiple analysis updates, gradually capturing a reasonable posterior’s geometry [chada_iterative_2020, blomker_continuous_2021], which facilitates information quantification and outer-loop design optimization.
@@ 2.4 AD-EKI for A Differentiable Estimation of Expected Information Gain
$\tilde{D}_{\text{KL}}$ represents the actual information gain provided by some measurement data from the true system. In the context of experimental design, those measurement data are often unavailable a priori before an experiment is designed. Therefore, the expected information gain (EIG) is employed to design an experiment based on the modeled system and prior knowledge without requiring the actual data from the experiments. As expressed in Eq. 5, EIG integrates out the dependency on data $\mathbf{y}$ in the KL divergence $D_{\text{KL}}$ by the prior predictive distribution $p(\mathbf{y}|\mathbf{d})$. The differentiation formulae of $\tilde{D}_{\text{KL}}$ presented in Section 2.3 remains valid, except for that data $\mathbf{y}$ is predicted by the model when EIG is employed as the utility function of BED.
Specifically, the distribution $p(\mathbf{y}|\mathbf{d})$ of the predicted data $\mathbf{y}$ is induced by the prior distribution $p(\bm{\theta})$ of the parameter $\bm{\theta}$, which is propagated through the PDE solution process under a given design $\mathbf{d}$ and subsequently transformed by the observation model with observing noise $\bm{\eta}\sim\mathcal{N}(\mathbf{0},\Gamma)$: $\{p(\bm{\theta}),\mathcal{N}(\mathbf{0},\Gamma)\}\mapsto p(\mathbf{y}|\mathbf{d})$. While directly evaluating this predictive distribution $p(\mathbf{y}|\mathbf{d})$ and calculating $\tilde{D}_{\text{KL}}(\mathbf{d,\mathbf{y}^{(m)}})$ for each sampled $\mathbf{y}^{(m)}\sim p(\mathbf{y}|\mathbf{d})$ is possible, this sampling process is not differentiable and reparameterization trick [kingma_auto-encoding_2013] is required.
@@
$\mathbb{E}_{\mathbf{y(\mathbf{d})}\sim p(\mathbf{y}|\mathbf{d})}[\tilde{D}_{\text{KL}}(\mathbf{d},\mathbf{y})]=\mathbb{E}_{\bm{\theta}\sim p(\bm{\theta}),\mathbf{\mathbf{\eta}}\sim p(\eta)}\{\tilde{D}_{\text{KL}}[\mathbf{d},\mathbf{y}(\bm{\theta},\mathbf{\eta})]\}\approx\frac{1}{M}\sum_{m=1}^{M}\tilde{D}_{\text{KL}}[\mathbf{d},\mathbf{y}(\bm{\theta}^{m},\mathbf{\eta}^{m})],$ (14)
where each $\tilde{D}_{\text{KL}}[\mathbf{d},\mathbf{y}(\bm{\theta}^{m},\mathbf{\eta}^{m})]$ is evaluated by Eq. (9) in this work. Consequently, differentiating EIG with respect to design $\mathbf{d}$ can be estimated as:
$\frac{\mathrm{d}}{\mathrm{d}\mathbf{d}}\mathbb{E}_{\mathbf{y(\mathbf{d})}}[\tilde{D}_{\text{KL}}(\mathbf{d},\mathbf{y})]$ $\approx\frac{\mathrm{d}}{\mathrm{d}\mathbf{d}}\left[\frac{1}{M}\sum_{m=1}^{M}\tilde{D}_{\text{KL}}[\mathbf{d},\mathbf{y}(\bm{\theta}^{m},\mathbf{\eta}^{m})]\right]$ (15)
$=\frac{1}{M}\sum_{m=1}^{M}\frac{\mathrm{d}}{\mathrm{d}\mathbf{d}}\tilde{D}_{\text{KL}}[\mathbf{d},\mathbf{y}(\bm{\theta}^{m},\mathbf{\eta}^{m})],$
where the interchange of the order of differentiation and integration in the second equal is valid with mild assumptions that $\tilde{D}_{\text{KL}}$ and $\partial\tilde{D}_{\text{KL}}/\partial\mathbf{d}$ are continuous and an integrable function bounds the partial derivative term.
For a single sample of $\mathbf{y}^{m}$, the computational cost of AD-EKI remains nearly unchanged. As the calculation of $\tilde{D}_{\text{KL}}[\mathbf{d},\mathbf{y}(\bm{\theta}^{m},\mathbf{\eta}^{m})]$ for each sample is independent, the total computational complexity for $M$ samples scales linearly as $\mathcal{O}(MJ)$. It is also worth noting that sequentially performing AD-EKI for each sample does not change the peak memory cost of $\mathcal{O}(J)$. The estimation of EIG in Eq. (14) differs from a nested Monte Carlo method, which is commonly used in BED to approximate the double integral in EIG. The primary distinction between this formulation of Eq. (14) and the nested Monte Carlo method lies in the estimation of $\tilde{D}_{\text{KL}}[\mathbf{d},\mathbf{y}(\bm{\theta}^{m},\mathbf{\eta}^{m})]$, which is approximated by the ensemble Kalman inversion in this work. A key advantage is that a small ensemble size is often sufficient for ensemble Kalman methods, even if dealing with a high-dimensional problem, e.g., only an order of $10^{2}$ samples are typically required to handle parameter spaces with dimensions as high as $10^{8}$ in numerical weather prediction, with various techniques developed for a robust covariance estimation [tong_localization_2023, vishny_high-dimensional_2024].
@@ 2.5 A Hybrid Framework for BED with Model Discrepancy
In the context of BED with model discrepancy, parameterizing the discrepancy in a data-driven way as Eq. (2) leads to two distinct sets of unknowns: the physical parameters $\bm{\theta}_{\mathcal{G}}$ and the network parameters $\bm{\theta}_{\textrm{NN}}$. Although performing BED on the joint parameter space $\{\bm{\theta}_{\mathcal{G}},\bm{\theta}_{\textrm{NN}}\}$ is a natural first thought, incorporating network parameters significantly increases the dimensionality of the parameter space, posing a major challenge for full Bayesian methods due to computational intractability. It is worth noting that some recent works [wu2023large, wu2023fast, go2025sequential, go2025accurate] have been developed for Bayesian experimental design with high-dimensional unknown parameters, which rely on building an efficient and generalizable parameter-to-observable map and thus are not directly applicable for the modeled system in Eq. (2). As an efficient alternative to full Bayesian methods, ensemble Kalman methods introduce Gaussian assumptions, are capable of handling high-dimensional parameter spaces, and demonstrate robust performance on nonlinear problems. However, the true distribution of physical parameters may be strongly non-Gaussian. Therefore, we propose a hybrid strategy that treats physical parameters with a full Bayesian approach, while dealing with neural network parameters via AD-EKI and gradient-based optimization. More specifically, we decompose the BED problem for $\{\bm{\theta}_{\mathcal{G}},\bm{\theta}_{\textrm{NN}}\}$ into two sub-problems. First, we determine the optimal design for the physical parameters $\bm{\theta}_{\mathcal{G}}$ using standard BED techniques and update the belief distribution through Bayesian inference. Next, we determine the optimal experiment design for the calibration of neural network parameters $\bm{\theta}_{\textrm{NN}}$ by maximizing the ensemble-based utility function enabled by AD-EKI in Section 2.4. With the data from the optimal design, we then update $\bm{\theta}_{\textrm{NN}}$ via a standard gradient-based optimization.
@@
$\mathbf{d}^{*}_{\text{NN}}=\operatorname*{arg\,max}_{\mathbf{d}\in\mathcal{D}}\mathbb{E}[\tilde{D}_{\text{KL}}(p(\bm{\theta}_{\text{NN}}|\mathbf{y};\bm{\theta}_{\mathcal{G}}^{*})||p(\bm{\theta}_{\text{NN}}))],$ (19)
where $\tilde{D}_{\text{KL}}$ denotes the KL divergence estimated by the AD-EKI introduced in Section 2.3. For each stage of sequential BED with model correction, we assume a Gaussian prior over the neural network coefficients centered at their current values. With the optimal design $\mathbf{d}^{*}_{\text{NN}}$ being identified for the neural-network-based model discrepancy, the corresponding data $\mathbf{y}_{\text{NN}}$ is then used to update the network coefficients by a standard gradient-based optimization method. In practice, we leverage the entire accumulated dataset $\mathrm{Y}_{\text{NN}}^{(i-1)}=\{\mathbf{y}_{\text{NN}}^{1},\mathbf{y}_{\text{NN}}^{2},\cdots,\mathbf{y}_{\text{NN}}^{(i-1)}\}$ from previous stages in sequential BED, together with the data $\mathbf{y}_{\text{NN}}^{(i)}$ at the stage $i$ for the training of the neural-network-based model discrepancy.
A detailed algorithm for sequential BED with model discrepancy is summarized in Algorithm 1.
Algorithm 1 Hybrid BED framework with Model Discrepancy via AD-EKI
for i=1,2,,Ni=1,2,...,N do $\triangleright$ Iterating sBED stage
  BED for physical parameter
  $\mathbf{d}_{\mathcal{G}}\leftarrow\arg\max_{\mathbf{d}\in\mathcal{D}}\mathbb{E}[U(\bm{\theta}_{\mathcal{G}},\mathbf{y},\mathbf{d};\bm{\theta}_{\textrm{NN}})]$ $\triangleright$ Traditional BED with $\bm{\theta}_{\textrm{NN}}$ fixed
  $\mathbf{y}_{\mathcal{G}}\leftarrow\mathbf{u}(\mathbf{d}_{\mathcal{G}})+\bm{\epsilon}$ $\triangleright$ Measurement at optimal design
  $\bm{\theta}_{\mathcal{G}}\leftarrow\bm{\theta}_{\mathcal{G}}|\mathbf{y}_{\mathcal{G}}$ $\triangleright$ Bayesian update
  BED for network parameter
  $\bm{\theta}_{\mathcal{G}}^{*}\leftarrow\operatorname*{arg\,max}_{\bm{\theta}_{\mathcal{G}}}p(\bm{\theta}_{\mathcal{G}};\bm{\theta}_{\text{NN}})$ $\triangleright$ MAP
  $\mathbf{d}_{\textrm{NN}}\leftarrow\arg\max_{\mathbf{d}\in\mathcal{D}}\mathbb{E}[\tilde{D}_{\textrm{KL}}(p(\bm{\theta}_{\textrm{NN}}|\mathbf{y})||p(\bm{\theta}_{\textrm{NN}});\bm{\theta}_{\mathcal{G}}^{*})]$ $\triangleright$ AD-EKI with $\bm{\theta}_{\mathcal{G}}^{*}$ fixed
  $\mathbf{y}_{\textrm{NN}}\leftarrow\mathbf{u}(\mathbf{d}_{\textrm{NN}})+\bm{\epsilon}$ $\triangleright$ Measurement at optimal design
  $\bm{\theta}_{\textrm{NN}}\leftarrow\max_{\bm{\theta}_{\text{NN}}}L(\bm{\theta}_{\text{NN}};\bm{\theta}_{\mathcal{G}}^{*},\mathbf{y}_{\text{NN}},\mathbf{d}_{\text{NN}})$ $\triangleright$ Gradient-based optimization
  $\bm{\theta}_{\mathcal{G}}\leftarrow\bm{\theta}_{\mathcal{G}}|\mathbf{y}_{\mathcal{G}};\bm{\theta}_{\text{NN}}$ $\triangleright$ Bayesian update conditioned on updated $\bm{\theta}_{\text{NN}}$
end for